# Predicting Bad Goods Risk Scores with ARIMA Time Series: A Novel Risk Assessment Approach

### v1

@@ I INTRODUCTION
In contemporary industrial ecosystems, the identification and prevention of defective or substandard products—commonly referred to as “bad goods”—pose significant challenges. Bad goods, such as manufacturing defects or spoiled perishables like Organic Beer-G 1 Liter, lead to substantial financial losses, reputational damage, and inefficiencies across supply chains. Traditional quality control methods, including statistical process control and manual inspections, often fall short in addressing the dynamic, multifaceted nature of these issues, particularly in large-scale operations. The proliferation of big data and advanced computational techniques has positioned predictive analytics as a transformative tool to preemptively detect and mitigate such risks [1].
This research proposes a predictive analytics framework that combines Time Series ARIMA models with risk assessment to forecast the occurrence of bad goods while prioritizing interventions for supply chain stakeholders. Time series data, capturing sequential patterns of sales, returns, capacity, and other supply chain variables, forms the foundation for modeling temporal trends. The ARIMA model, known for its ability to handle linear and stationary time series, is employed to predict bad goods risks, addressing the limitations of conventional forecasting methods like Exponential Smoothing [2]. Additionally, a risk assessment component, inspired by methodologies such as Failure Mode and Effects Analysis (FMEA), quantifies the severity and probability of quality failures, enabling effective resource allocation [3].
The motivation for this study arises from the shortcomings of reactive quality management systems and the need for a proactive, data-driven solution to minimize bad goods in industries like beverage production and retail. The objectives are threefold: (1) to develop an accurate predictive model for bad goods using Time Series ARIMA, (2) to integrate risk assessment for actionable insights, and (3) to evaluate the framework’s performance against established benchmarks. This work has profound implications for manufacturing, retail, and logistics sectors, where reducing defective goods can yield significant cost savings and improve customer satisfaction.
@@ II BASIC CONCEPTS
Quality Control: Quality control in supply chains ensures product integrity, minimizing defects and returns. For perishable goods like beer, factors such as freshness, shelf life, and capacity play a pivotal role in determining bad goods risks. This study leverages quality control principles to guide ARIMA predictions and risk assessments, ensuring alignment with operational realities and industry standards [4].
@@ III RELATED WORKS
The application of predictive analytics to quality control and defect detection in supply chains has garnered increasing attention in recent years. Traditional statistical methods, such as Autoregressive Integrated Moving Average (ARIMA) models, have been extensively utilized for time series forecasting in industrial contexts [1]. While effective for linear and stationary data, ARIMA models face challenges with non-linear, high-dimensional, or noisy datasets common in modern supply chains, prompting exploration of alternative approaches [2]. For instance, machine learning techniques like Support Vector Machines (SVM) and Random Forests have been applied for anomaly detection and defect prediction in manufacturing processes, but these methods often struggle to model long-term temporal dependencies critical for bad goods prediction [4].
@@ VI-G Implications for Beer-G Product
The statistical analysis indicates that Beer-G experiences significant volatility in retailer capacity, bought quantity, and return quantity, with weak autocorrelation suggesting short-term randomness rather than long-term trends. The strong correlation between purchases and returns highlights potential quality or satisfaction issues that warrant further investigation, such as product defects or marketing misalignments. The right-skewed distribution of ‘rate_of_return‘ and ‘return_qty‘, alongside the multimodal distributions of ‘bought_qty‘ and ‘retailer_capacity‘, suggest variability in customer behavior and operational constraints. The forecasted stabilization of both retailer capacity and rate of return in 2025–2026, combined with the historical decline in capacity after mid-2024, indicates a potential market adjustment or reduced demand. Retailers and manufacturers of Beer-G should focus on improving product quality, optimizing inventory management, addressing return drivers, and monitoring capacity utilization to enhance performance and customer satisfaction.
@@ VI-H Interpretation of Bad Goods Risk Scores
1 Smith et al.[7] ARIMA Regression SupplyChainDB
2 Johnson et al.[8] Exponential Smoothing Decision Trees RetailData
3 Lee et al.[9] LSTM Deep Learning SCMetrics
4 Patel et al.[10] Holt-Winters Statistical Models ChainAnalytics
@@ 5 Proposed Work ARIMA Time Series ARIMA Collected
These recommendations aim to balance demand, capacity, and freshness to minimize the risk of bad goods, particularly for months classified as Medium or High Risk, ensuring product quality, reducing returns, and optimizing supply chain operations.
VII COMPARISON OF OUR WORK WITH PRESENT STATE-OF-THE-ART TECHNIQUES
Time series analysis has emerged as a powerful tool for predicting bad goods risks in supply chain management, offering the ability to forecast trends and assess risks based on historical data patterns. Several researchers have explored the application of time series models, such as ARIMA, Exponential Smoothing, and LSTM (Long Short-Term Memory), for risk prediction in supply chains. In this section, we compare our work with some of the existing state-of-the-art techniques, as shown in Table II.
Our proposed approach leverages Time Series ARIMA integrated with a comprehensive risk assessment framework to predict bad goods risk scores, distinguishing it from other methods that often focus solely on forecasting without risk scoring or use different statistical or machine learning techniques. Unlike Smith et al.[7] and Johnson et al.[8], which rely on simpler regression and decision tree models without risk assessment, our method incorporates both prediction and risk scoring, enhancing decision-making for supply chain stakeholders. Compared to Lee et al.[9], who use deep learning (LSTM) for both prediction and risk scoring, our ARIMA-based approach is computationally lighter and more interpretable, while achieving comparable performance. Patel et al.[10] use Holt-Winters for forecasting but lack risk assessment, limiting their applicability for bad goods management. Our work’s integration of ARIMA with risk scoring, applied to a diverse and sizable dataset, positions it as a novel and effective solution for bad goods risk prediction in the supply chain.

### v2

@@ I INTRODUCTION
In modern industrial systems, detecting and preventing defective or substandard products—termed “bad goods”—such as manufacturing flaws or spoiled items like Organic Beer-G 1 Liter, remains a critical challenge. These defects result in financial losses, reputational harm, and supply chain inefficiencies. Traditional approaches like statistical process control and manual inspections struggle to address the complexity of large-scale operations [1]. The advent of big data and advanced analytics has elevated predictive methods as a key strategy for preempting such risks [2].
This study introduces a predictive analytics framework that integrates Time Series ARIMA models with risk assessment to forecast bad goods occurrences and guide supply chain interventions. Leveraging time series data—encompassing sales, returns, and capacity—ARIMA effectively models temporal patterns, outperforming simpler methods like Exponential Smoothing in capturing linear trends [2]. A risk assessment layer, inspired by Failure Mode and Effects Analysis (FMEA), quantifies failure likelihood and impact, enhancing decision-making [3].
Motivated by the limitations of reactive quality management, this research aims to deliver a proactive, data-driven solution for industries such as beverage production and retail. The objectives are to: (1) build a precise ARIMA-based predictive model, (2) incorporate risk assessment for practical insights, and (3) benchmark its efficacy against existing standards. By reducing bad goods, this approach promises cost savings and improved customer satisfaction across manufacturing, retail, and logistics sectors.
@@ II BASIC CONCEPTS
Quality Control: Quality control in supply chains ensures product integrity, minimizing defects and returns. For perishable goods like beer, factors such as freshness, shelf life, and capacity play a pivotal role in determining bad goods risks. This study leverages quality control principles to guide ARIMA predictions and risk assessments, ensuring alignment with operational realities and industry standards.
@@ III RELATED WORKS
The application of predictive analytics to quality control and defect detection in supply chains has garnered increasing attention in recent years. Traditional statistical methods, such as Autoregressive Integrated Moving Average (ARIMA) models, have been extensively utilized for time series forecasting in industrial contexts [1]. While effective for linear and stationary data, ARIMA models face challenges with non-linear, high-dimensional, or noisy datasets common in modern supply chains, prompting exploration of alternative approaches [2]. For instance, machine learning techniques like Support Vector Machines (SVM) and Random Forests have been applied for anomaly detection and defect prediction in manufacturing processes, but these methods often struggle to model long-term temporal dependencies critical for bad goods prediction.
@@ VI-G Implications for Beer-G Product
The statistical analysis shows Beer-G exhibits notable volatility in retailer capacity, purchases, and returns, with weak autocorrelation indicating short-term randomness rather than sustained trends. A strong purchase-return correlation suggests potential quality or satisfaction issues, meriting further exploration into defects or marketing discrepancies.
@@ VI-H Interpretation of Bad Goods Risk Scores
These recommendations aim to balance demand, capacity, and freshness to minimize the risk of bad goods, particularly for months classified as Medium or High Risk, ensure product quality, reduce returns, and optimize supply chain operations.
VII COMPARISON OF OUR WORK WITH PRESENT STATE-OF-THE-ART TECHNIQUES
Time series analysis has become a vital tool for predicting risks associated with defective or ”bad” goods in supply chain management, enabling the forecasting of trends and the evaluation of risks based on historical data. Researchers have employed various time series models—such as ARIMA, Exponential Smoothing, Holt-Winters, and LSTM (Long Short-Term Memory)—to address supply-chain risk prediction
@@ VII COMPARISON OF OUR WORK WITH PRESENT STATE-OF-THE-ART TECHNIQUES
1 Hyndman et al. [6] Exponential Smoothing Statistical Simulated
2 Taylor [7] Holt-Winters Statistical Retail Sales
3 Zhang et al. [8] LSTM Deep Learning Supply Chain Data
4 Billah et al. [9] ARIMA Statistical Economic Data
@@ 5 Proposed Work ARIMA Time Series ARIMA Collected
challenges. In this section, we compare our proposed approach with existing state-of-the-art techniques, as summarized in Table II.
Our method integrates Time Series ARIMA with a robust risk assessment framework to predict bad goods risks and assign risk scores, setting it apart from techniques that focus solely on forecasting or employ alternative statistical and machine learning approaches. For instance, Hyndman et al. [6] utilize Exponential Smoothing for forecasting but do not incorporate risk assessment, limiting its utility for proactive risk management. Similarly, Taylor [7] applies the Holt-Winters method to retail sales data for prediction, yet it lacks a risk scoring component. In contrast, Zhang et al. [8] leverage LSTM for both prediction and risk assessment in supply chain contexts, though their deep learning approach is computationally intensive and less interpretable than our ARIMA-based solution. Billah et al. [9] employ ARIMA for forecasting economic time series but do not extend it to risk scoring, unlike our work.
Our proposed approach, by combining ARIMA with risk scoring and applying it to a comprehensive collected dataset, offers a computationally efficient, interpretable, and effective solution for bad goods risk prediction in supply chains. This dual focus on prediction and risk assessment enhances decision-making capabilities for supply chain stakeholders, positioning our method as a valuable contribution to the field.